# Counterfactual Uncertainty Quantification of Factual Estimand of Efficacy from Before-and-After Treatment Repeated Measures Randomized Controlled Trials

### v3

@@ Abstract
The ideal estimand for comparing treatment $Rx$ with a control $C$ is the counterfactual efficacy $Rx:C$, the expected differential outcome between $Rx$ and $C$ if each patient were given both. One hundred years ago, Neyman (1923a) proved unbiased point estimation of counterfactual efficacy from designed factual experiments is achievable. But he left the determination of how much might the counterfactual variance of this estimate be smaller than the factual variance as an open challenge. This article shows counterfactual uncertainty quantification (CUQ), quantifying uncertainty for factual point estimates but in a counterfactual setting, is achievable for Randomized Controlled Trials (RCTs) with Before-and-After treatment Repeated Measures which are common in many therapeutic areas. We achieve CUQ whose variability is typically smaller than factual UQ by creating a new statistical modeling principle called ETZ.
We urge caution in using predictors with measurement error which violates standard regression assumption and can cause attenuation in estimating treatment effects. Fortunately, we prove that, for traditional medicine in general, and for targeted therapy with efficacy defined as averaged over the population, counterfactual point estimation is unbiased. However, for both Real Human and Digital Twins approaches, predicting treatment effect in subgroups may have attenuation bias.
Keywords: Variability decomposition, Randomized Controlled Trials, Before-and-After Repeated measures, Digital Twins, Counterfactual uncertainty quantification, ETZ modelling principle.
@@ 1 Motivation and scope of application
Most factual Randomized Controlled Trials (RCTs) are such that each patient is given either $Rx$ or $C$ but not both. However, counterfactual point estimation of counterfactual $Rx:C$ efficacy from factual studies is readily available, as random assignment of treatments to patients lets one prove that the point estimate of the differential $Rx$ and $C$ effects from modeling a factual RCT is unbiased for the counterfactual $Rx:C$ efficacy. The proof that designed experiments allow for unbiased counterfactual point estimation of the average treatment effect (ATE) from observable outcomes can be traced to Neyman (1923b), which has been edited and translated as Splawa-Neyman et al. (1990). Importantly, Neyman realized that the variance of a factual point estimate in standard practice is larger than what the variance of a counterfactual point estimate would be, and considered the problem of accurately assessing the latter “worthy of future study”. See page 473 of Rubin (1990) which commented on Neyman (1923a).
@@
To gain insight, we propose a new statistical modeling principle called ETZ which separates variability of patient outcomes at the milestone visit into three components: variability in patients’ initial severity of the disease ($Z$), variability in trajectories ($Traj$) of patients’ responses to treatments, and variability in the measurement errors ($E$). ETZ is non-standard in that $Z$ is not directly observable.
Our ETZ modeling principle in Section 3.5 shows using change-from-baseline instead of the milestone visit outcome reduces the uncertainty in efficacy assessment so long as the measurement error variability is not the main source of variability at the baseline visit (visit 1), a minimal requirement (for otherwise the outcome scale is of questionable usefulness). So the ETZ modeling principle validates current practice.
@@
A final insight from the ETZ modeling principle is as follows. Since $Z$ is unobservable, it needs to be estimated. While standard practice has been to estimate it using baseline, more generally it can be estimated using values of a set of biomarkers (with baseline as a special case). What we realized is that, with regression analysis assuming predictors are measured without error, using predictors measured with error causes a so-called attenuation in estimating treatment effects. Fortunately, after careful analysis, we discovered that counterfactual point estimation is always unbiased for traditional medicine. As long as efficacy is defined as averaged over the population, counterfactual point estimation is also unbiased for targeted therapies as shown in Section 4.5. However, all approaches (including our real human self-controlling as well as the “digital twins” approach) should be aware of the limitation that predicting treatment effect in subgroups may have bias, if one or more of the biomarker values are measured with error, as shown in Section 4.4.
@@ 2.1 Conceptualizing an ideal counterfactual efficacy
2.2 Conceptualizing an ideal personal $Z$ covariate
@@
Compared to the usual baseline predictor, the $Z$ component has two unique features: it is without measurement error, and it is not observable. As visualized in Figure 1, neither the intercept $Z$ nor the trajectory curves are directly observable, because there are measurement errors. Observed and potential outcomes are represented as dots and squares respectively in Figure 1.
We call ETZ a counterfactual modeling principle because $Z$ is not observable. Each person has a scalar $Z$ for their covariate $X$, on the $Y$ outcome scale. The ETZ modeling principle uses $Z$ to model biological variability among patients, conceptualizing initial intercept value $Z$ as predictive of outcome.
The ETZ modeling principle reduces the quantified uncertainty of estimated counterfactual efficacy at the population level by first using each patient as their own control (as in current practice), then further reduces it by calculating counterfactually under the ETZ modeling principle. We believe the resulting sample size reduction we achieve is at least as competitive as any other approaches, perhaps more so.
@@ 2.3 A factual Digital Twins model
Instead of conceptualizing patients as having different (unobservable) initial values $Z$, one can also account for patient variability by their having different observable biomarker values $(b_{1},\ldots,b_{p})$ such as age, baseline measurement, severity of disease, sex, etc. An algorithm $g$ can be trained to predict an $Rx$ patient’s digital twin’s $C$ outcome and a $C$ patient’s digital twin’s $Rx$ outcome based on $(b_{1},\ldots,b_{p})$. In what we call Digital Twin Counterfactual Estimation (DTCE), after calculating the difference between each patient’s factual and its digital twin’s counterfactual outcome, population $Rx:C$ efficacy is estimated by averaging them across the patients. DTCE can also potentially reduce uncertainty of estimated efficacy.
@@ 2.5 Independence as a starting point
Independence means that, a patient’s true severity of illness before they receives either $Rx$ or $C$ does not inform on whether that patient’s response trajectory is going to be above or below the profile $\mu$ (long run average of the trajectories). The concept of Independence is disconnected with whether Z is prognostic of Change or not. The concept of baseline being prognostic will be defined in Section 4.1. In addition, note that independence between $Z$ and $Traj$ is within each treatment arm, unrelated to whether baseline and treatment should have an interaction term in the model or not.
If patient level data is available, then one way to check independence is to see how baseline measurement is predictive of change-from-baseline. We take an AbbVie Alzheimer’s disease (AD) data as an example. Schnell et al. (2017) described an Alzheimer Disease study from AbbVie which compared three doses of a new treatment (doses 1, 2, and 3) with a negative control (dose 0, a placebo). There was an active control (dose 4) in the study as well. Outcome measure was ADAS-cog11, and duration of this Phase 2 study was 24 weeks. For the AbbVie AD data, the R-square regressing change-from-baseline (negative of the “improve” variable in the data set) on baseline (variable “severity” in the data set) allowing different profiles for the placebo and the 4 doses was 0.17. So it appears severity is not very predictive of trajectory.
A subtle concern is, regression modeling assumes predictors are measured without error, what if some of them have measurement error? Fortunately, as will be explained in Section 4.5, so long as efficacy is defined as “on average”, estimated $Rx:C$ efficacy is unbiased. However, that Section 4.4 also shows, in predicting efficacy in subgroups or at particular biomarker values, ignoring measurement error in predictors will lead to bias.
@@ 2.7 BAtRM for counterfactual inference
ETZ is a modeling principle applicable to any BAtRM data structure, not restricted to any specific analysis method. How to estimate $\sigma^{B}$, $\sigma^{Mile}$, and $\sigma^{Change}$ from factual data is not unique. For proof of concept, we show how they can be obtained from widely used Mixed Model Repeated Measures (MMRM) analyses introduced in 3.1, and then transformed 1-to-1 to variabilities in the ETZ modeling principle in Section 3.4.
@@ 3.1 Mixed Model Repeated Measures (MMRM)
can all be estimated.
@@ 3.3 Plausibility
Intuitively, variability at the first and milestone visits and of Change may contain all the information we need to separate the $E$, $Traj$, and $Z$ variability components. Baseline measurements at the beginning of the study (visit 1) have the random intercept $Z$ and measurement error $E$ components but no random trajectory component $Traj$ since patients have not been treated yet. Change, or change-from-baseline, measurement at the milestone visit minus measurement at the first visit, has the random trajectory $Traj$ and the measurement error $E$ components (two copies of $E$ in fact) but not the random intercept component since that has been subtracted off. Measurements at the end of the study (visit $m$) have all three random components. So, assuming variability of measurement errors is constant across the visits, it seems plausible that $Var(Y^{[1]}_{i})$, $Var(Y^{[m]}_{i})$, and $Var(Y^{[\mathrm{change}]}_{i})$ are sufficient for us to recover the the $E$, $Traj$, and $Z$ variability components.
@@ 3.4 Components of the ETZ decomposition
with the number of unknown ETZ variability components remaining five. Making the assumption $Var(E^{[1]}_{r})=Var(E^{[m]}_{r})$, reducing the number of unknowns to four, gives
@@
Note that, if $Z$ and $Traj$ are in fact positively correlated (say), then from (16), (18), and (3.4), estimates of $Var(Z_{r})$ and $Var(Traj^{Trt[m]}_{r})$ would be somewhat higher than they ought to be, while the estimate for $Var(E^{[1]}_{r})=Var(E^{[m]}_{r})$ would be somewhat lower than it ought to be.
@@ 3.5 Uncertainty reduced by counterfactual uncertainty quantification
We now show exactly how counterfactual uncertainty quantification (CUQ), the goal of our research, is achieved by the ETZ modeling principle, increasing the power of clinical trials, decreasing the sample size needed.
@@
Baselining very effectively reduces uncertainty
@@
We show below that baselining effectively replaces the patient-to-patient variability $Var(Z)$ by the measurement error variability $Var(E)$ and, so long as $Var(Z)>Var(E)$, baselining is beneficial.
@@ 3.6 Real data examples of CUQ
Figure 2: Counterfactual uncertainty quantification can reduce the quantified uncertainty in efficacy estimation. The pie charts on the left show the proportion of total variance reduced by baselining while the pie charts on the right show the additional proportion reduced by counterfactual (instead of factual) calculations. The top two pie charts analyze data from the AbbVie study, while the bottom two pie charts analyze data from the EXPEDITION3 study. Overall, counterfactual uncertainty quantification reduces 77.19% and 39.74% of the variances in the AbbVie study and the EXPEDITION3 study, respectively.
@@ 4.1 Genesis of baseline covariate adjustment
If the data contains substantial imbalance, then a real example in Hsu (1992) showed the marginal analysis implemented in the Means statement of SAS Proc GLM which does not adjust for the prognostic factor as a covariate can give very biased $Rx:C$ efficacy estimates.
@@ 4.2 From traditional medicine to targeted therapy
Suppose in practice, for patients $r$ and $s$, $B_{r}=Z_{r}+\prescript{{}_{B}}{}{E}^{[1]}_{r}$ and $B_{s}=Z_{s}+\prescript{{}_{B}}{}{E}^{[1]}_{s}$ with measurement errors are used instead of their (not observable) $Z$ values for prediction and ordinary regression is applied to the models (27) and (28).
@@ 4.3 Noisy predictor causes attenuation
We show below the presence of measurement error $\prescript{{}_{B}}{}{E}$ or $E$ in using baseline or any other biomarker $B$ score as a surrogate for the true patient effect $Z$ causes attenuation in estimating the regression coefficient of $B$, which in turn can cause bias in estimating $Rx:C$ efficacy. While future ETZ research might mitigate this bias, for now we show that estimated counterfactual efficacy is unbiased so long as efficacy is defined as “on average”.
@@ 4.4 Possible bias caused by attenuation
visit 1 (baseline) 64.58 8.04
visit 7 (milestone visit) 135.39 11.64
Change-from-baseline 92.37 9.61
Intercept Z 53.802 7.335
Trajectory 70.809 8.415
Measurement Error 10.778 3.283
Table 4: Variance and standard deviation (SD) of iADL of visit 1, visit $m$, and change-from-baseline of the EXPEDITION3 study. Variances of Intercept, Trajectory, and Measurement Error are computed by ETZ transformation.
Figure 6: Black arrows in the plot show attenuation bias in estimated $C$ effect, due to measurement error in biomarker score $B$ as a surrogate of the true $Z$. Treatments are indicated by color while shape differentiates $B$ from $Z$. Red represents control $C$ while teal represents treatment $Rx$. Red round points are $C$ versus $B$ measurements, with the red dashed line being the fitted $C$ versus $B$ regression line. Red triangle points are $C$ versus $Z$ measurements, with the red solid line being the fitted $C$ versus $Z$ regression line. Teal triangle points are $Rx$ versus $Z$ measurements, with the teal solid line being the fitted $Rx$ versus $Z$ regression line.
Figure 7: The direction of bias in $Rx:C$ efficacy estimation depends on the target value of outcome $C$. Efficacy is overestimated when target $C$ value is below average, and underestimated when target $C$ value is above average. For the red points, the (horizontal) abscissa (coordinate) is outcome $C$ as predicted by $B$, and the (vertical) ordinate (coordinate) is outcome $Rx$ as predicted by $Z$. The red line is the fitted line of $Rx$ as predicted by true $Z$ regressed on $C$ as predicted by $B$. For the teal points, both the abscissa outcome $C$ and the ordinate outcome $Rx$ are as predicted by true $Z$. The teal line is the fitted line of $Rx$ as predicted by true $Z$ regressed on $C$ as predicted by true $Z$. For easier visual perception, the $Rx$ and $C$ outcome values plotted have been partially denoised using values predicted by $Z$ after we reduce the standard deviation of $Z$ from the original 7.34 to 3.
Using baseline or, more generally, a biomarker score $B$ to predict the outcome is tempting because $Z$ is not observable. However, as Figures 6 and 7 show, one should be aware that predicting the expected outcome at a $B$ value off-center may incur bias if $B$ has measurement error.

### v4

@@ Abstract
This article quantifies the uncertainty reduction achievable for counterfactual estimand, and cautions against potential bias when the estimand uses Digital Twins.
Posed by Neyman (1923a) who showed unbiased point estimation from designed factual experiments is possible, counterfactual uncertainty quantification (CUQ) remained an open challenge for about one hundred years. The $Rx:C$ counterfactual efficacy we focus on is the ideal estimand for comparing treatment $Rx$ with control $C$, the expected outcome differential if each patient received both $Rx$ and $C$.
Enabled by our new statistical modeling principle called ETZ, we show CUQ is achievable in Randomized Controlled Trials (RCTs) with Before-and-After Repeated Measures, common in many therapeutic areas. The CUQ we are able to achieve typically has lower variability than factual UQ.
We caution against using predictors with measurement error, which violates regression assumptions and can cause attenuation bias in estimating treatment effects. For traditional medicine and population-averaged targeted therapy, counterfactual point estimation remains unbiased. However, in both Real Human and Digital Twin approaches, estimating effects in subgroups may suffer attenuation bias.
Keywords: Variability decomposition, Randomized Controlled Trials, Digital Twins, Counterfactual Uncertainty Quantification, ETZ modelling principle, Before-and-After Repeated Measures.
@@ 1 Motivation and scope of application
Most factual Randomized Controlled Trials (RCTs) are such that each patient is given either $Rx$ or $C$ but not both. However, counterfactual point estimation of counterfactual $Rx:C$ efficacy from factual studies is readily available, as random assignment of treatments to patients lets one prove that the point estimate of the differential $Rx$ and $C$ effects from modeling (average treatment effect (ATE)) a factual RCT is unbiased for the counterfactual $Rx:C$ efficacy.
The proof that designed experiments allow for unbiased counterfactual point estimation of the average treatment effect (ATE) from observable outcomes can be traced to Neyman (1923b), which has been edited and translated as Splawa-Neyman et al. (1990).
Importantly, how to quantify the uncertainty in estimating a counterfactual point estimate is a problem raised one hundred years ago but has remained unsolved. Neyman considered this problem “worthy of future study”. Neyman is unable to assess the variance of a counterfactual point estimate precisely, although he realized that the variance of a factual point estimate in standard practice is larger than what the variance of a counterfactual point estimate would be. See page 473 of Rubin (1990) which commented on Neyman (1923a).
@@
To gain insight, we propose a new statistical modeling principle called ETZ which separates variability of patient outcomes at the milestone visit into three components: variability in severity (stage of the disease that patients are initially at) $Z$, variability in trajectories $Traj$ of patients’ responses to treatments, and variability in the measurement errors $E$.
Our ETZ modeling principle in Section 4.5 shows using change-from-baseline instead of the milestone visit outcome reduces the uncertainty in efficacy assessment provided that the measurement error variability is not the main source of variability at the baseline visit (visit 1), a minimal requirement (for otherwise the outcome scale is of questionable usefulness). So the ETZ modeling principle validates current practice.
@@
A final insight from the ETZ modeling principle is that, since $Z$ is unobservable, it must be estimated. While baseline is often used, $Z$ can more generally be estimated from a set of biomarkers. However, when predictors are measured with error, standard regression analysis leads to attenuation of treatment effect estimates. Fortunately, we found that counterfactual point estimation remains unbiased for traditional medicine. For targeted therapies, it is also unbiased, as long as efficacy is defined as population-averaged (see Section 5.5). That said, all methods including our self-controlled human design and digital twins—should note that subgroup-level effect estimation may be biased if biomarkers are measured with error (see Section 5.4).
@@ 2.1 Conceptualizing an ideal counterfactual efficacy
3 Conceptualizing an ideal personal $Z$ covariate
@@
Compared to the usual baseline predictor, the $Z$ component has two unique features: it is unobservable and without measurement error. As visualized in Figure 1, neither the intercept $Z$ nor the trajectory curves are directly observable, because there are measurement errors. Observed and potential outcomes are represented as dots and squares respectively in Figure 1.
We call ETZ a counterfactual modeling principle because $Z$ is not observable. Each individual is assigned a scalar $Z$ for their covariate $X$, on the $Y$ outcome scale. The ETZ modeling principle uses $Z$ to model biological variability among patients, conceptualizing initial intercept value $Z$ as predictive of outcome.
The ETZ modeling principle reduces the quantified uncertainty of estimated counterfactual efficacy at the population level by first using each patient as their own control (as in current practice), then further reduces it by calculating counterfactually under the ETZ modeling principle.
@@ 3.1 A factual Digital Twins model
Instead of conceptualizing patients as having different (unobservable) initial values $Z$, one can also account for patient variability by their having different observable biomarker values $(b_{1},\ldots,b_{p})$ such as age, baseline or other measurement of disease severity, sex, etc. An algorithm $g$ can be trained to predict an $Rx$ patient’s digital twin’s $C$ outcome and a $C$ patient’s digital twin’s $Rx$ outcome based on $(b_{1},\ldots,b_{p})$. In what we call Digital Twin Counterfactual Estimation (DTCE), after calculating the difference between each patient’s factual and its digital twin’s counterfactual outcome, population $Rx:C$ efficacy is estimated by averaging them across the patients. See National Academies of Sciences, Engineering, and Medicine (2024), for example, for such thinking. DTCE can also potentially reduce uncertainty of estimated efficacy.
@@ 3.3 Independence as a starting point
Independence means that, the true stage of the disease that a patient is at before they receives either $Rx$ or $C$ does not inform on whether that patient’s response trajectory is going to be above or below the profile $\mu$ (long run average of the trajectories). By the Monotone Class Theorem, it is sufficient to check whether $P\{B|A\}=P\{B\}$ for events $A$ and $B$ of the form $B=\{improve>b\}$ and $A=\{severity>a\}$. That is, within each treatment arm, whether knowledge of severity being greater than 25 (say) informs on whether a patient would improve by more than 10 points (say). (Note that independence between the random effects $Z$ and $Traj$ within each treatment arm is unrelated to whether the fixed effect response profiles are parallel or not.)
If patient-level data is available, then one way to check independence is to see how baseline measurement is predictive of change-from-baseline. We take an AbbVie Alzheimer’s disease (AD) data as an example. Schnell et al. (2017) described an Alzheimer Disease study from AbbVie which compared three doses of a new treatment (doses 1, 2, and 3) with a negative control (dose 0, a placebo). There was an active control (dose 4) in the study as well. Outcome measure was ADAS-cog11, and duration of this Phase 2 study was 24 weeks. For the AbbVie AD data, the R-square regressing change-from-baseline (negative of the “improve” variable in the data set) on baseline (variable “severity” in the data set) allowing different profiles for the placebo and the 4 doses is 0.17. So it appears severity is not very predictive of trajectory.
A subtle concern is, regression modeling assumes predictors are measured without error, what if some of them have measurement error? Fortunately, as will be explained in Section 5.5, as long as efficacy is defined as “on average”, estimated $Rx:C$ efficacy is unbiased. However, that Section 5.4 also shows, in predicting efficacy in subgroups or at particular biomarker values, ignoring measurement error in predictors will lead to bias.
@@ 3.5 BAtRM for counterfactual inference
ETZ is a modeling principle applicable to any BAtRM data structure, not restricted to any specific analysis method. How to estimate $\sigma^{B}$, $\sigma^{Mile}$, and $\sigma^{Change}$ from factual data is not unique. For proof of concept, we show how they can be obtained from widely used Mixed Model Repeated Measures (MMRM) analyses introduced in Section 4.1, and then transformed 1-to-1 to variabilities in the ETZ modeling principle in Section 4.4.
@@ 4.1 Mixed Model Repeated Measures (MMRM)
[…]
The standard MMRM analysis from which we obtain these estimates assumes the missing data mechanism is MAR (Missing At Random). Strategies for dealing with other missing mechanisms are given in Chapter 19 of Mallinckrodt et al. (2019).
@@ 4.3 Plausibility
Intuitively, variability at the first and milestone visits and of Change may contain all the information we need to separate the $E$, $Traj$, and $Z$ variability components. Baseline measurements at the beginning of the study (visit 1) have the random intercept $Z$ and measurement error $E$ components but no random trajectory component $Traj$ since patients have not been treated yet. Change, or change-from-baseline, measurement at the milestone visit minus measurement at the first visit, has the random trajectory $Traj$ and the measurement error $E$ components (two copies of $E$ in fact) but not the random intercept component since that has been subtracted off. Measurements at the end of the study (visit $m$) have all three random components. So, assuming variability of measurement errors is constant across the visits, it seems plausible that $Var(Y^{[1]}_{i})$, $Var(Y^{[m]}_{i})$, and $Var(Y^{[\mathrm{change}]}_{i})$ are sufficient for us to recover the $E$, $Traj$, and $Z$ variability components.
@@ 4.4 Components of the ETZ decomposition
with the number of unknown ETZ variability components remaining five. (Note that since estimates of the three variances in (16) are routinely included as summary statistics in reporting RCT outcomes, $Cov(Y^{Trt[m]}_{r},Y^{Trt[1]}_{r})$ in (17) can be estimated without access to patient level data.) Making the assumption $Var(E^{[1]}_{r})=Var(E^{[m]}_{r})$, reducing the number of unknowns to four, gives
@@
Figure 2 examines limitations of current ETZ modeling with the independence assumption. One potential issue is, if $Z$ and $Traj$ are in fact positively correlated, then from (17), (18), and (4.4), estimates of $Var(Z_{r})$ and $Var(Traj^{Trt[m]}_{r})$ would be higher than they ought to be, while the estimate for $Var(E^{[1]}_{r})=Var(E^{[m]}_{r})$ would be lower than it ought to be. Then we might miss the signal that $Var(E)>Var(Z)$, in which case study result is unreliable because within-patient measurement variability dominates between-patient variability. The right panel of Figure 2 shows how one can make a plot to assess how large the correlation needs to be for this signal to be missed.
Figure 2: Ratio SD($E$)/SD($Z$) of standard deviations across different clinical endpoints (ADAS-Cog, iADL, and MMSE) as functions of the correlation between $Z$ and $Traj$, using EXPEDITION3 as an example. The left panel shows if $Z$ and $Traj$ are sufficiently negatively correlated, then ETZ modeling can result in $Var(E)<0$ clearly indicating the independence assumption does not hold. The right panel shows the extent to which positive correlation between $Z$ and $Traj$ can cause ETZ modeling with the independence assumption to miss the signal that $Var(E)>Var(Z)$. Purple dashed line indicates when SD ($E$) reaches 80% of SD ($Z$), while orange dashed line indicates SD ($E$) = SD ($Z$).
Another potential issue is, if $Z$ and $Traj$ are sufficiently negatively correlated, then as shown in the left panel of Figure 2, ETZ modeling can result in $Var(E)<0$ which is nonsensical. But we do not expect negative correlation to be common in practice, as that corresponds to patients at late stage of the disease having more favorable trajectories.
@@ 4.5 Uncertainty reduced by counterfactual uncertainty quantification
We now show exactly how Counterfactual Uncertainty Quantification (CUQ), the goal of our research, is achieved by the ETZ modeling principle.
@@
Baselining significantly reduces uncertainty
@@
We show below that baselining effectively replaces the patient-to-patient variability $Var(Z)$ by the measurement error variability $Var(E)$ and, provided that $Var(Z)>Var(E)$, baselining is beneficial.
@@ 4.6 Real data examples of CUQ
Figure 3: Counterfactual uncertainty quantification can reduce the quantified uncertainty in efficacy estimation. The pie charts on the left show the proportion of total variance reduced by baselining while the pie charts on the right show the additional proportion reduced by counterfactual (instead of factual) calculations. The top two pie charts analyze data from the AbbVie study, while the bottom two pie charts analyze data from the EXPEDITION3 study. Overall, Counterfactual Uncertainty Quantification reduces 77.19% and 39.74% of the variances in the AbbVie study and the EXPEDITION3 study, respectively. The detailed calculation has been demonstrated in Appendix B.
@@ 5.1 Genesis of baseline covariate adjustment
If the data contains a substantial imbalance, then a real example in Hsu (1992) showed the marginal analysis implemented in the Means statement of SAS Proc GLM which does not adjust for the prognostic factor as a covariate can give very biased $Rx:C$ efficacy estimates.
@@ 5.2 From traditional medicine to targeted therapy
In practice, for patients $r$ and $s$, $B_{r}=Z_{r}+\prescript{{}_{B}}{}{E}^{[1]}_{r}$ and $B_{s}=Z_{s}+\prescript{{}_{B}}{}{E}^{[1]}_{s}$ with measurement errors are used instead of their (not observable) $Z$ values for prediction and ordinary regression is applied to the models (27) and (28). Burman et al. (2024) cites PROCOVA as an example of Digital Twin, deriving a prognostic score from historical patient-level baseline data and using it as a covariate in ANCOVA modeling to boost trial efficiency. Appendix D of Schuler et al. (2022) presents the predictors PROCOVA used to generate the prognostic score. Measurements of a number of the predictors, such as ADAS comprehension, CDR home and hobbies, and MMSE language, are subject to measurement error.
@@ 5.3 Noisy predictor causes attenuation
We show below the presence of measurement error $\prescript{{}_{B}}{}{E}$ or $E$ in using baseline or any other biomarker $B$ score as a surrogate for the true patient effect $Z$ causes attenuation in estimating the regression coefficient of $B$, which in turn can cause bias in estimating $Rx:C$ efficacy. While future ETZ research might mitigate this bias, for now we show that estimated counterfactual efficacy is unbiased provided that efficacy is defined as “on average”.
@@ 5.4 Possible bias caused by attenuation
Visit 1 (baseline, $Y^{[1]}$) 64.58 8.04
Visit 7 (milestone visit, $Y^{[m]}$) 135.39 11.64
Change-from-baseline ($Y^{[change]}$) 92.37 9.61
Intercept ($Z$) 53.80 7.34
Trajectory ($Traj$) 70.81 8.42
Measurement Error ($E$) 10.78 3.28
Table 4: Variance and standard deviation (SD) of iADL at Visits 1 and $m$, and change-from-baseline of the EXPEDITION3 study are as reported in Honig et al. (2018). Vriances of the Intercept, Trajectory and Measurement Error are computed by the ETZ transformation.
Figure 7: This plot shows attenuation when $B$ is a surrogate of the true $Z$ measured with error. The slope regressing $Y$ on $B$ tends to be closer to zero, compared to the true slope regressing $Y$ on $Z$. Black arrows in the plot show attenuation bias in estimated $C$ effect, due to measurement error in biomarker score $B$ as a surrogate of the true $Z$. Treatments are indicated by color while shape differentiates $B$ from $Z$. Red represents control $C$ while teal represents treatment $Rx$. Red round points are $C$ versus $B$ measurements, with the red dashed line being the fitted $C$ versus $B$ regression line. red dashed line are $C$ versus $Z$ measurements, with the red solid line being the fitted $C$ versus $Z$ regression line. Teal triangle points are $Rx$ versus $Z$ measurements, with the teal solid line being the fitted $Rx$ versus $Z$ regression line.
Figure 8: This plot shows how bias in $Rx$ versus $C$ efficacy estimation depends on the target value of outcome $C$. The orange points represent outcome $Rx$ (vertical axis) as predicted by true $Z$ plotted against outcome $C$ (horizontal axis) as predicted by true $Z$, with the orange line being the prediction line. The blue points represent outcome $Rx$ (vertical axis) as predicted by true $Z$ plotted against outcome $C$ (horizontal axis) as predicted by $B(=Z+E)$, with the blue line being their fitted line. $Rx$ versus $C$ efficacy, the average difference of response given $Rx$ and response given $C$, is biased upward when target $C$ value is below average (with the blue points trending above the orange points), downward when target $C$ value is above average (with blue points trending below orange points).
Using baseline or, more generally, a biomarker score $B$ to predict the outcome is tempting because $Z$ is not observable. However, as Figures 7 and 8 show, one should be aware that predicting the expected outcome at a $B$ value off-center may incur bias if $B$ has measurement error.